It costs a manufacturer $7.95 to produce sunglasses that sell for $25.99. How many pairs of sunglasses must the manufacturer sell to make a profit of $17,138?
pairs of sunglasses ?

Answer by ikleyn(52781)

.

    25.99x = 7.95x + 17138  dollars

    25.99x - 7.95x = 17138

        18.04x     = 17138

             x     = 17138/18.04 = 950.    ANSWER

Solved.